# Do We Need All the Synthetic Data? Towards Targeted Synthetic Image Augmentation via Diffusion Models

### v1

@@ 4.2 Synthetic Data Augmentation Amplifies Features But not Noise
Recall that $D$ is the original dataset with $|D|=N$. We assume exactly $(1-\alpha)N\in\mathbb{Z}$ samples have only ${\bm{v}}_{d}$, and let $D_{U}$, $D_{G}$ be the modified datasets via upsampling and generation with factor $k$ and new size $N_{new}=\alpha N+k(1-\alpha)N$. For $D_{U}$, the replicated noises $\{\bm{\xi}_{i}:i=\alpha N+1,\dots,N\}$ introduce a dependence. Additionally, for $D_{G}$, the generative model will have its own noises $\bm{\gamma}_{i}$ (potentially higher than original) for the new data, and we assume the noises are i.i.d., orthogonal to features, and independent from feature noise.
@@
With similar notations, let $\bm{\Phi}_{G}$, $\bm{\Phi}_{U}$ denote the multi-sets  [^1] Due to identical noises in $D_{U}$. The following ${\mathcal{I}}_{j,+}^{U,(t)}$ and ${\mathcal{I}}_{j,-}^{U,(t)}$ are also multi-sets to keep their sizes consistent with the generation counterparts, meaning that we count all the replicated noises. of all the noises for $D_{G}$, $D_{U}$ respectively. We define ${\mathcal{I}}_{j,+}^{G,(t)}=\{\bm{\phi}_{i}\in\bm{\Phi}_{G}:i\in[N_{new}],\ \ \text{sgn}\langle{\bm{w}}_{j}^{(t)},\bm{\phi}_{i}\rangle=\text{sgn}(y_{i})\}$, ${\mathcal{I}}_{j,-}^{G,(t)}$, ${\mathcal{I}}_{j,+}^{U,(t)}$, ${\mathcal{I}}_{j,-}^{U,(t)}$ in a similar fashion for $D_{G}$ and $D_{U}$ as before.
@@
Now we are ready to present the following theorem, which suggests that early in training, for feature learning, upsampling and generation contribute similarly, but upsampling accelerates noise learning for noises in ${\mathcal{I}}_{j,+}^{U,(t)}$, while synthetic generation does not.
@@
Let $\nabla_{{\bm{w}}_{j}^{{(t)}}}^{U}({\mathcal{L}}({\bm{W}}^{(t)}))$ and $\nabla_{{\bm{w}}_{j}^{{(t)}}}^{G}({\mathcal{L}}({\bm{W}}^{(t)}))$ denote the full gradients w.r.t the $j$-th neuron for upsampling and generation at iteration $t$ respectively. With sufficiently large $N$, $J$, sufficiently small learning rate $\eta$ and noises, starting with the same weights early in training, we expect that for one gradient update,
@@
    $|\langle\nabla_{{\bm{w}}_{j}}^{U}{\mathcal{L}}({\bm{W}}^{(t)}),\bm{\xi}_{i}\rangle|=\begin{cases}\frac{3}{N_{new}}l_{i}^{(t)}\langle{\bm{w}}_{j}^{(t)},\bm{\xi}_{i}\rangle^{2}\|\bm{\xi}_{i}\|^{2},&i=1,\dots,\alpha N\\ \frac{3k}{N_{new}}l_{i}^{(t)}\langle{\bm{w}}_{j}^{(t)},\bm{\xi}_{i}\rangle^{2}\|\bm{\xi}_{i}\|^{2},&i=\alpha N+1,\dots,N\end{cases}.$
A special case of this theorem is that with the same initializations ${\bm{W}}^{(0)}\sim\mathcal{N}(0,\sigma_{0}^{2})$, informally, we expect $\textbf{NoiseAlign}({\mathcal{I}}_{j,+}^{G,(t)},{\bm{w}}_{j}^{(t)})<\textbf{NoiseAlign}({\mathcal{I}}_{j,+}^{U,(t)},{\bm{w}}_{j}^{(t)})$ for all the early iterations $0\leq t\leq T$, if the generation noise is sufficiently small in the sense that in expectation,
$\langle{\bm{w}}_{j}^{(0)},\bm{\gamma}\rangle^{2}\|\bm{\gamma}\|^{2}<(k+1)\langle{\bm{w}}_{j}^{(0)},\bm{\xi}\rangle^{2}\|\bm{\xi}\|^{2}\quad\text{for some data noise $\bm{\xi}$ and generation noise $\bm{\gamma}$.}$
Theorem 4.2 captures one important observation in practice that not all the noises are created equal. Additionally, we note that for generation, the update rules learn each noise in a similar fashion, but for upsampling, we still only have $N$ unique noises, and the training dynamics deviates more from SAM for the extra learning on the upsampled ones. Consequently, choosing $k>2$ increasingly pushes the model toward overfitting particular noises—explaining the empirically observed performance drop of upsampling at larger $k$ for Useful.
Remark. Intuitively, for a diffusion model, as long as the noises in the synthetic data $\|\bm{\gamma}_{i}\|^{2}$ are not too large, noise learning under generation is more diffused due to its independence and operates more similarly to SAM, whereas upsampling amplifies learning on the duplicated noise, potentially elevating it to a new feature and increasing the NoiseAlign metric.
@@ 4.3 Superior Convergence of Synthetic Data Augmentation over Upsampling
Next, we compare the convergence of mini-batch Stochastic Gradient Descent (SGD)—which is used in practice—on synthetically augmented and upsampled data and confirm the superior convergence of synthetic data augmentation. Consider training the above CNN using SGD with batch size B:
${\bm{w}}_{j}^{(t+1)}={\bm{w}}_{j}^{(t)}-\eta\frac{1}{B}\sum_{i=1}^{B}\nabla_{i,{\bm{w}}_{j}^{{(t)}}}{\mathcal{L}}({\bm{W}}^{(t)}).$
The convergence rate of mini-batch SGD is inversely proportional to the batch size [17]. The following theorem shows that upsampling inflates the variance of mini-batch gradients and consequently slows down convergence. Note that it does not directly compare the relative magnitudes of the variances, since $\sigma_{G}(k)$, $\sigma_{U}(k)$ depend on datasets. Instead, it quantifies the sources of extra variance:
@@
From Theorem 4.3, we see that all the variance for generation comes from the per-example variance, potentially getting larger when synthetic images diversify the dataset. However, upsampling induces one extra term that results from repeated noises and unnecessarily inflates the variance due to dependence within the dataset. This is empirically justified in our ablation studies in Section 5.4.
@@
As long as the generation noise is small enough, i.e., $\sigma_{G}^{2}(k)<\sigma_{U}^{2}(k)$, convergence of mini-batch SGD on synthetically augmented data is faster than upsampled data.
@@ 4.4 Generating Faithful Synthetic Images via Diffusion Models
Synthetic image generation with diffusion models [36, 22, 32, 42] involves a forward process to iteratively add noise to the images, followed by a reverse process to learn to denoise the images. Specifically, the forward process progressively adds noise to the data $x_{0}$ over $T$ steps, with each step modeled as a Gaussian transition: $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I})$, where $\beta_{t}$ controls the noise added at each step. The reverse process inverts the forward process, learns to denoise the data, with the goal of recovering the original data $x_{0}$ from a noisy sample $x_{T}$: $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t))$, where mean $\mu_{\theta}$ is conditioned on the sample at the previous time step and variance $\Sigma_{\theta}$ follows a fixed schedule.
To ensure generating images that are faithful to real data, we use the real images as guidance to generate synthetic images. Specifically, while using the class name (e.g., “a photo of a dog") as the text prompt, we also incorporated the original real samples as guidance. More formally, instead of sampling a a pure noisy image $x_{T}\sim\mathcal{N}(0,I)$ as the initialization of the reverse path, we add noise to a reference (real) image $x_{0}^{ref}$ such that the noise level corresponds to a certain time-step $t_{*}$. Then we begin denoising from time-step $t_{*}$, using an open-source text-to-image model, e.g. GLIDE [31], to iteratively predict a less noisy image $x_{t-1}(t=T,T-1,\cdots,1)$ using the given text prompt $l$ and the noisy latent image $x_{t}$ as inputs. This technique enables produce synthetic images that are similar, yet distinct, from the original examples, and has been successfully used for synthetic image generation for few-shot learning [19].
The pseudocode of our method is illustrated in Alg. 1 in Appendix E. Our algorithm is similar to Useful except that in Step 3, we replace the real images by synthetic images generated as above.
@@ 5 Experiment
5.1 Settings
Base training datasets. We use common benchmark datasets for image classification including CIFAR10, CIFAR100 [27], and Tiny-ImageNet [28] for image classification.
Augmented training datasets. We train different models on: (1) Original: The original training datasets without any modifications. (2) Useful: Augmented dataset with upsampled real images that are not learned in early training. (3) Our Method: Replace the upsampled samples in Usefulwith their corresponding synthetic images. For $k>2$, we use generated images at different diffusion denoising steps, instead of generating from scratch.
Training details. We train ResNet18 on all datasets following the setting of [2]. The models are trained for 200 epochs with a batch size of 128. We use SGD with the momentum parameter 0.9 and set weight decay to 0.0005. We also fix $\rho=0.1$ for SAM in all experiments. We use a linear learning rate schedule starting at 0.1 and decay by a factor of 10 once at epoch 100 and again at epoch 150. For each setup, we perform training using two different optimization methods: SGD and SAM.
Hyperparameters. For Useful [30], we adopt their default hyper-parameters for separating epochs and set the upsampling factor $k$ to 2 as it yields the best performance. For our method, the upsampling factor $k$ is set to $5$ for CIFAR10 and CIFAR100 while that of TinyImageNet is set to $4$. For generating synthetic images using GLIDE, we use guidance scale of 3 and run denoising for 100 steps, saving generated images every 10 steps.
More details about the experimental settings can be found in Appendix F.1.
@@ 5.2 Our Method is Effective across Datasets and Architectures
Different datasets. Figure 1 clearly shows that our method significantly reduces the test classification error compared to both the Original and Useful methods across all datasets, namely CIFAR10, CIFAR100, and Tiny-ImageNet. For Tiny ImageNet, our method yields an improvement of more than 2% regardless of the optimizer choice. The superior performance of our method compared to Useful is well algined with our theoretical results in Section 4.3. Notably, SGD with our method outperforms SAM on CIFAR100 and TinyImageNet. This clearly confirms the effectiveness of our approach.
@@
Figure 1: Test classification error of ResNet18 on CIFAR10, CIFAR100 and TinyImageNet. For Useful, the upsampling factor $k$ is always set to 2 while for Ours, $k$ is set to 5, 5, and 4 for CIFAR10, CIFAR100, and Tiny-ImageNet, respectively. Our method improves both SGD and SAM. Notably, it enables SGD to outperform SAM on CIFAR100 and TinyImageNet.
@@
Different model architectures. To further evaluate the generalization of our approach, we conduct experiments on multiple model architectures using CIFAR10 as the base dataset. Specifically, we apply our method to CNNs (VGG19, DenseNet121) and Transformers (ViT-S). Figure 2 presents the test classification error for different architectures. The results demonstrate that our method achieves consistently lower classification error than both the Original and USEFUL methods across all architectures, under both SGD and SAM optimization settings. These findings further confirm the robustness and effectiveness of our approach across different network structures.
@@
Figure 2: Test classification error of VGG19, DenseNet121, and ViT-S on CIFAR10. For Useful, the upsampling factor $k$ is set to 2 while for Ours, $k$ is set to 5.
@@
Figure 3: (left & middle) Comparison between different synthetic image augmentation strategies when training ResNet18 on CIFAR10 and CIFAR100. The upsampling factor $k$ is set to 2 for Syn-rand and Ours, resulting in only 30% and 40% additional examples compared to 100% of Syn-all. (right) Our method with $k=5$ can be stacked with TrivialAugment (TA) to further boosts the performance when training ResNet18 on CIFAR10, achieving (to our knowledge) SOTA test classification error.
@@
Figure 4: (left) The effect of the upsampling factor $k$ on the performance of training ResNet18 with SGD on the augmented CIFAR10 dataset produced by our method and Useful. Red points indicate the optimal choice of $k$ for each method. (middle) Comparison between synthetic CIFAR10 images generated from random noises and real images when training ResNet18 on the augmented set with $k=2$. For reference, we include the performance of the original training set. (right) The effect of the number of denoising steps on the performance of training ResNet18 on the augmented CIFAR10 dataset produced by our method with $k=2$.
5.3 Do We Need All the Synthetic Data?
To answer the question “Do We Need All the Synthetic Data?", we mix all the corresponding generated images with the training data at a 1:1 ratio, doubling the training set size, denoted as Syn-all. In this case, we set $k=2$ in our method, resulting in increase of approximately 30% and 40% of the total training data in CIFAR10 and CIFAR100, respectively. This means that the computational cost of generating these samples is much lower than that of Syn-all, which requires 100% of the training data to be augmented with generated images. Consequently, the generation time for our method is reduced to 0.3x and 0.4x that of Syn-all, making it more efficient. In addition, we consider a baseline (Syn-rand) where random images are augmented with their corresponding synthetic ones. In terms of the test classification error, our method outperforms both Syn-rand (same cost) and Syn-all (higher cost) as shown in Figure 3. This highlights the effectiveness of our targeted augmentation compared to random augmentation or simply augmenting all the images.

### v2

@@ 4.2 Synthetic Data Augmentation To Amplify Features But not Noise
Recall that $D$ is the original dataset with $|D|=N$. We assume exactly $(1-\alpha)N\in\mathbb{Z}$ samples have only ${\bm{v}}_{d}$, and let $D_{U}$, $D_{G}$ be the modified datasets via upsampling and generation with factor $k$ and new size $N_{new}=\alpha N+k(1-\alpha)N$. For $D_{U}$, the replicated noises $\{\bm{\xi}_{i}:i=\alpha N+1,\dots,N\}$ introduce a dependence. Additionally, for $D_{G}$, the generative model will have its own noises $\bm{\gamma}_{i}$ (potentially higher) for synthetic data, and we assume the noises are i.i.d., orthogonal to features, and independent from feature noise from some distribution ${\mathcal{D}}_{\bm{\gamma}}$.
@@
With similar notations, let $\bm{\Phi}_{G}$, $\bm{\Phi}_{U}$ denote the multi-sets  [^1] Due to identical noises in $D_{U}$. The following ${\mathcal{I}}_{j,+}^{U,(t)}$ and ${\mathcal{I}}_{j,-}^{U,(t)}$ are also defined via multi-sets to keep their sizes consistent with the generation counterparts, meaning that we count all the replicated noises. of all the noises for $D_{G}$, $D_{U}$ respectively. We define ${\mathcal{I}}_{j,+}^{G,(t)}=\{\bm{\phi}_{i}\in\bm{\Phi}_{G}:i\in[N_{new}],\ \ \text{sgn}\langle{\bm{w}}_{j}^{(t)},\bm{\phi}_{i}\rangle=\text{sgn}(y_{i})\}$, ${\mathcal{I}}_{j,-}^{G,(t)}$, ${\mathcal{I}}_{j,+}^{U,(t)}$, ${\mathcal{I}}_{j,-}^{U,(t)}$ in a similar fashion for $D_{G}$ and $D_{U}$ as before.
@@
Next, we show that early in training, upsampling and generation contribute similarly to feature learning, but upsampling accelerates learning noises in ${\mathcal{I}}_{j,+}^{U,(t)}$, while synthetic generation does not.
@@
Let $\nabla_{{\bm{w}}_{j}^{{(t)}}}^{U}({\mathcal{L}}({\bm{W}}^{(t)}))$ and $\nabla_{{\bm{w}}_{j}^{{(t)}}}^{G}({\mathcal{L}}({\bm{W}}^{(t)}))$ denote the full gradients w.r.t the $j$-th neuron for upsampling and generation at iteration $t$ respectively. When dropping the index $i$, the term is treated as a random variable. Then for one gradient update,
@@
    $|\langle\nabla_{{\bm{w}}_{j}}^{U}{\mathcal{L}}({\bm{W}}^{(t)}),\bm{\xi}_{i}\rangle|=\begin{cases}\frac{3}{N_{new}}l_{i}(t)\langle{\bm{w}}_{j}^{(t)},\bm{\xi}_{i}\rangle^{2}\|\bm{\xi}_{i}\|^{2},&i=1,\dots,\alpha N\\ \frac{3k}{N_{new}}l_{i}(t)\langle{\bm{w}}_{j}^{(t)},\bm{\xi}_{i}\rangle^{2}\|\bm{\xi}_{i}\|^{2},&i=\alpha N+1,\dots,N\end{cases}.$
A special case of this theorem is that with the same initializations ${\bm{W}}^{(0)}\sim\mathcal{N}(0,\sigma_{0}^{2})$, for all the early iterations $0\leq t\leq T$, if the synthetic noises are sufficiently small in the sense that
$\mathbb{E}\left[l_{\bm{\gamma}}(t)\langle{\bm{w}}_{j}^{(t)},\bm{\gamma}\rangle^{2}\|\bm{\gamma}\|^{2}\right]$ $<(k+1)\mathbb{E}\left[l_{\bm{\xi}}(t)\langle{\bm{w}}_{j}^{(t)},\bm{\xi}\rangle^{2}\|\bm{\xi}\|^{2}\right],$
where $l_{\bm{\gamma}}(t)$, $l_{\bm{\xi}}(t)$ denote the logits as random variables under the corresponding noise at iteration $t$,
$\text{then noises are overfitted less: }\quad\mathbb{E}\left[\textbf{NoiseAlign}({\mathcal{I}}_{j,+}^{G,(t)},{\bm{w}}_{j}^{(t)})\right]<\mathbb{E}\left[\textbf{NoiseAlign}({\mathcal{I}}_{j,+}^{U,(t)},{\bm{w}}_{j}^{(t)})\right].$
Theorem 4.2 suggests that synthetically generating faithful images that preserve features in real data with independent noise prevents noise overfitting in expectation. In particular, this metric of generation noise does not need to be strictly less than that of data noise during early training; instead, there exists a tolerance factor up to $k+1$.
Remark. For a diffusion model, as long as the noises in the synthetic data $\|\bm{\gamma}_{i}\|^{2}$ are not too large, noise learning under generation is more diffused due to its independence and operates more similarly to SAM, whereas upsampling amplifies learning on the duplicated noise, potentially elevating it to a new feature and increasing the NoiseAlign metric.
@@ 4.3 Superior Convergence of Synthetic Data Augmentation over Upsampling
Next, we show the superior convergence of mini-batch Stochastic Gradient Descent (SGD)—which is used in practice—on synthetically augmented compared to upsampled data. Consider training the above CNN using SGD with batch size B: ${\bm{w}}_{j}^{(t+1)}={\bm{w}}_{j}^{(t)}-\eta\frac{1}{B}\sum_{i=1}^{B}\nabla_{i,{\bm{w}}_{j}^{{(t)}}}{\mathcal{L}}({\bm{W}}^{(t)})$. The convergence rate of mini-batch SGD is inversely proportional to the batch size (Ghadimi & Lan, 2013). The following theorem shows that upsampling inflates the variance of mini-batch gradients and thus slows down convergence. Note that it does not directly compare the relative magnitudes of the variances, since $\sigma_{G}(k)$, $\sigma_{U}(k)$ depend on datasets. Instead, it quantifies the sources of extra variance:
@@
From Theorem 4.3, we see that all the variance for generation solely comes from the per-example variance, potentially getting larger when synthetic images diversify the dataset. However, upsampling induces one extra term that results from repeated noises and unnecessarily inflates the variance due to dependence within the dataset. This is empirically justified in our ablation studies in Section 5.2.
@@
As long as the generation noise is small enough, i.e., $\sigma_{G}^{2}(k)\leq\sigma_{U}^{2}(k)$, convergence of mini-batch SGD on synthetically augmented data is faster than upsampled data.
@@ 4.4 Generating Faithful Synthetic Images via Diffusion Models
Synthetic image generation with diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Nichol & Dhariwal, 2021; Yang et al., 2023) involves a forward process to iteratively add noise to the images, followed by a reverse process to learn to denoise the images. Specifically, the forward process progressively adds noise to the data $x_{0}$ over $T$ steps, with each step modeled as a Gaussian transition: $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I})$, where $\beta_{t}$ controls the noise added at each step. The reverse process inverts the forward process, learns to denoise the data, with the goal of recovering the original data $x_{0}$ from a noisy sample $x_{T}$: $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t))$, where mean $\mu_{\theta}$ is conditioned on the sample at the previous time step and variance $\Sigma_{\theta}$ follows a fixed schedule.
To ensure generating images that are faithful to real data, we use the real images as guidance to generate synthetic images. Specifically, while using the class name (e.g., “a photo of a dog”) as the text prompt, we also incorporated the original real samples as guidance. More formally, instead of sampling a a pure noisy image $x_{T}\sim\mathcal{N}(0,I)$ as the initialization of the reverse path, we add noise to a reference (real) image $x_{0}^{ref}$ such that the noise level corresponds to a certain time-step $t_{*}$. Then we begin denoising from time-step $t_{*}$, using an open-source text-to-image model, e.g. GLIDE (Nichol et al., 2021), to iteratively predict a less noisy image $x_{t-1}(t=T,T-1,\cdots,1)$ using the given text prompt $l$ and the noisy latent image $x_{t}$ as inputs. This technique enables produce synthetic images that are similar, yet distinct, from the original examples, and has been successfully used for synthetic image generation for few-shot learning (He et al., 2023). Our pseudocode is illustrated in Appendix E Alg. 1.
@@ 5 Experiment
Base training datasets. We use common benchmark datasets for image classification including CIFAR10, CIFAR100 (Krizhevsky et al., 2009), Tiny-ImageNet (Le & Yang, 2015), Flowers-102 (Nilsback & Zisserman, 2008), Aircraft (Maji et al., 2013), and Stanford Cars (Krause et al., 2013).
Augmented training datasets. We train different models on: (1) Original: The original training datasets without any modifications. (2) Upsample: Augmented dataset with upsampled real images that are not learned in early training. (3) Our Method: Replace the upsampled samples in Upsamplewith their corresponding synthetic images. For $k>2$, we use generated images at different diffusion denoising steps, instead of generating from scratch.
Training details. We train ResNet18 on all datasets following the setting of (Andriushchenko & Flammarion, 2022). The models are trained for 200 epochs with a batch size of 128. We use SGD with the momentum parameter 0.9 and set weight decay to 0.0005. We also fix $\rho=0.1$ for SAM. We use a linear learning rate schedule starting at 0.1 and decay by a factor of 10 once at epoch 100 and again at epoch 150. For each setup, we train with both SGD and SAM.
Hyperparameters. For Upsample, we set the upsampling factor $k$ to 2 as it yields the best performance. For our method, we use $k=5$ for CIFAR10 and CIFAR100 and $k=4$ for TinyImageNet. For generating synthetic images using GLIDE, we use guidance scale of 3 and run denoising for 100 steps, saving generated images every 10 steps. More details are in Appendix F.1.
@@ 5.1 Our Method is Effective across Datasets and Architectures
Different datasets. Figure 2 clearly shows that our method significantly reduces the test classification error compared to both the Original and Upsample methods across all datasets, namely CIFAR10, CIFAR100, and Tiny-ImageNet. For Tiny ImageNet, our method yields an improvement of 2.8% when training with SAM. The superior performance of our method compared to Upsample is well aligned with our theoretical results in Section 4.3. Notably, SGD with our method outperforms SAM on CIFAR100 and TinyImageNet. This clearly confirms the effectiveness of our approach.
@@
Figure 2: Test classification error of ResNet18 on CIFAR10, CIFAR100 and TinyImageNet. For Upsample, we use a factor of $k=2$, as higher $k$ harms the performance. In contrast for our method (Ours), $k=5,5,4$ for CIFAR10, CIFAR100, and Tiny-ImageNet, respectively. Our method improves both SGD and SAM. Notably, it enables SGD to outperform SAM on CIFAR100 and TinyImageNet.
@@
Different model architectures. To further evaluate the generalization of our approach, we conduct experiments on multiple model architectures using CIFAR10 as the base dataset. Specifically, we apply our method to CNNs (VGG19, DenseNet121) and Transformers (ViT-S). Figure 3 presents the test classification error for different architectures. The results demonstrate that our method achieves consistently lower classification error than both the Original and Upsample methods across all architectures, under both SGD and SAM optimization settings. Moreover, when applied to state-of-the-art architectures such as ConvNeXt (Liu et al., 2022) and Swin Transformer (Liu et al., 2021), our method still outperforms the baselines by a substantial margin, as reported in Table 2. These findings confirm the effectiveness of our approach across different model structures.
Transfer learning and stacking with other synthetic image generations. We evaluate our method in a transfer learning setting, where we fine-tune a ResNet18 pre-trained on ImageNet-1K on the Flowers-102, Aircraft, and Stanford Cars datasets. Table 2 compares our method with DiffuseMix (Islam et al., 2024a) which is a SOTA data augmentation method. It enhances diversity of synthetic images by blending partial natural images with images generated via InstructPix2Pix (Brooks et al., 2023) diffusion model. Applying our method to augment slow-learnable images with DiffuseMix significantly outperform no-augmentation and augmenting entire dataset with DiffuseMix.
@@
Figure 3: Test classification error of VGG19, DenseNet121, and ViT-S on CIFAR10. For Upsample, we use a factor of k=2k=2—as higher $k$ hurst the performance—while for Ours, we use $k=5$.
Table 1: Test error of ConvNeXt-T and Swin-T on CIFAR-10 using SGD and k=2k\!=\!2.
Method ConvNeXtT SwinT
Original 37.33 $\pm$ 3.12 16.10 $\pm$ 0.19
Upsample 34.16 $\pm$ 2.47 14.93 $\pm$ 0.07
Ours 27.40 $\pm$ 1.99 14.57 $\pm$ 0.10
Table 2: Test error of pre-trained ResNet18 on Flowers-102, Aircraft, and Stanford Cars datasets.
Method Flowers-102 Aircraft Stanford Cars
Original 8.55 $\pm$ 0.19 26.02 $\pm$ 0.27 15.45 $\pm$ 0.02
DiffuseMix 8.92 $\pm$ 0.15 25.65 $\pm$ 0.20 15.19 $\pm$ 0.06
Ours+DifMx 8.08 $\pm$ 0.16 25.12 $\pm$ 0.35 14.96 $\pm$ 0.07
Do We Need All the Synthetic Data? To answer this question, we compare synthetically augmenting all examples (Syn-all)–which doubles the training set size–with our method with k=2k=2–which results in increase of approximately 30% and 40% of the total training data–in CIFAR10 and CIFAR100. Notably, the generation time for our method is reduced to 0.3x and 0.4x that of Syn-all, making it more efficient. In addition, we consider a baseline (Syn-rand) where random images are augmented with their corresponding synthetic ones. Figure 4 shows that our method has a much lower test classification error compared to Syn-rand (same cost) and Syn-all (higher cost). This highlights the effectiveness of our targeted data augmentation.
@@
Figure 4: (left & middle) Comparison between different synthetic image augmentation strategies when training ResNet18 on CIFAR10 and CIFAR100. For Syn-rand and Ours, we use $k=2$ resulting in only 30% and 40% additional examples compared to 100% of Syn-all. (right) Our method with $k=5$ can be stacked with TrivialAugment (TA) to further boosts the performance when training ResNet18 on CIFAR10, achieving (to our knowledge) SOTA test classification error.
@@
Figure 5: Training ResNet18 on CIFAR10. (left) The effect of amplification factor $k$ on test error for upsampling vs generation. Red points indicate the optimal choice of $k$. $k>2$ hurts upsampling but boosts generation. (middle) Generating synthetic CIFAR10 images from real images outperform starting from random noise. (right) Effect of the number of denoising steps on the performance with $k=2$.